# Deterministic Discrete Denoising

### v2

@@ Abstract
We propose a deterministic denoising algorithm for discrete-state diffusion models based on Markov chains. The generative reverse process is derandomized by introducing a variant of the herding algorithm with weakly chaotic dynamics, which induces deterministic discrete state transitions. Our approach is a direct replacement for the stochastic denoising process, requiring neither retraining nor continuous state embeddings. We demonstrate consistent improvements in both efficiency and sample quality on text and image generation tasks. Thus, this simple derandomization approach is expected to enhance the significance of discrete diffusion in generative modeling. Furthermore, our results reveal that deterministic reverse processes, well established in continuous diffusion, can also be effective in discrete state spaces.
@@ 1 Introduction
Diffusion models have recently achieved remarkable success in generating realistic image and audio data. Their forward process progressively corrupts data with noise, and the models are trained to denoise by inverting the corruption process. Once trained, diffusion models can generate realistic samples from pure noise by iteratively applying the denoising procedure. For continuous data such as images and speech, the reverse denoising process is often implemented as a deterministic algorithm, which exhibits more directed dynamics and enables efficient generation with fewer denoising steps compared to stochastic denoising, as demonstrated by DDIM (Song et al., 2021).
@@
Diffusion models for discrete data are also an important research direction, since many data of interest are inherently discrete, including text, graphs, genomes, and chemical structures. In discrete-state diffusion models, the forward and reverse processes are typically formulated as probabilistic Markov chains on a discrete sample space (Austin et al., 2021). If the reverse denoising process could be made deterministic, improved performance would be expected, as in the case of continuous models. However, in discrete spaces, naive derandomization as a deterministic mapping aggregates discrete trajectories and reduces sample diversity, which is essential for generative modeling.
@@
One promising approach is to embed the discrete sample space into a continuous space and apply continuous diffusion models (Hoogeboom et al., 2021; Chen et al., 2023; Sahoo et al., 2025). The main advantage of this approach is that it allows the direct use of well-established techniques for continuous diffusion, including deterministic denoising. Nevertheless, denoising in the continuous space requires retraining the model. Moreover, it remains unclear whether discrete data must be embedded into a continuous space for derandomization.
@@
In this paper, we propose to directly derandomize the reverse denoising process on discrete sample spaces. The key idea is to introduce a variant of the herding algorithm (Welling, 2009; Welling & Chen, 2010), which associates auxiliary continuous weight variables with each discrete variable in the sample space. The reverse denoising updates of both discrete and continuous variables are deterministic, and randomness arises solely from their initial states. Since the proposed algorithm derandomizes the state transitions of Markov chains, it serves as a drop-in replacement for the stochastic reverse process without retraining discrete diffusion models. This relation is analogous to DDIM (Song et al., 2021), which generates samples deterministically using a model trained for DDPM (Ho et al., 2020). In our experiments on text and image generation, we demonstrate consistent improvements in both efficiency and sample quality. These results reveal that deterministic reverse processes, well established in continuous diffusion, can also be effective in discrete state spaces.
@@
In the evaluation, we adopt UDLM (Schiff et al., 2025), a state-of-the-art discrete-state generative model based on the uniform diffusion process. Our aim in this paper is not to propose a new discrete diffusion model, but to demonstrate how the herding-based algorithm improves the generative process of existing ones. We use the official training, sampling, and evaluation codes of UDLM, replacing only the reverse denoising procedure. This directly shows that our method works as a drop-in replacement without retraining. Moreover, the uniform diffusion process is particularly suitable for our approach, since its reverse process performs iterative refinement through successive Markov state transitions, making it a natural setting to examine whether derandomizing these transitions can improve efficiency and sample quality.
In contrast, masked diffusion has been widely studied as a discrete diffusion model, but its reverse process allows each token to transition only once from the masked state and remain fixed thereafter, fundamentally limiting its dynamics. A related line of work is the deterministic reverse algorithm proposed in Chen et al. (2024), which performs deterministic state transitions in discrete sample spaces. This approach is effective for masked diffusion, where the transition time is limited. Our work complements this result by showing that deterministic denoising is also beneficial for uniform diffusion models, where state transitions continue throughout the reverse process.
Overall, we present a simple yet effective derandomization approach to directly enhance the generative process of existing discrete-state diffusion models, thus strengthening the significance of discrete diffusion in generative modeling.
2 Background
@@ 2.1 Discrete Diffusion Models
A diffusion process in a discrete state space can be viewed as a sequence of discrete state transitions that gradually corrupt the original data. This process can be formulated as a discrete-time Markov chain over categorical variables, giving rise to a reverse-time denoising Markov chain (Austin et al., 2021). This formulation can be naturally extended to state transitions in continuous time, using Markov jump processes parameterized by transition-rate matrices (Campbell et al., 2022; Sun et al., 2023). Foundational theoretical studies analyze approximation error and convergence properties of both discrete-time and continuous-time discrete diffusion models (Ren et al., 2025a; Zhang et al., 2025; Liang et al., 2025a; b). Below, we focus on a discrete-time formulation to outline how our approach can be applied to the reverse denoising process.
We consider generative models for discrete data represented as a sequence of $L$ categorical tokens. Each token takes one of $K$ categories and is represented as a $K$-dimensional one-hot vector $\bm{x}\in\mathcal{V}$, where $\mathcal{V}=\{(x_{1},x_{2},\dots,x_{K})^{\top}\in\{0,1\}^{K}\mid\sum_{k=1}^{K}x_{k}=1\}$. A sequence of $L$ tokens is denoted by $\bm{x}^{(1:L)}=(\bm{x}^{(1)},\bm{x}^{(2)},\dots,\bm{x}^{(L)})\in\mathcal{V}^{L}$.
@@
By iterating this Markov transition from pure noise $\bm{x}^{(1:L)}_{T}$ sampled from the stationary distribution, the reverse process generates a sample $\bm{x}^{(1:L)}_{0}$. Namely, each denoising step reduces to sampling from the categorical distribution $\mathrm{Cat}(\bm{x}_{t-1};\bm{p}_{t-1})$ with probability vector $\bm{p}_{t-1}=p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$. This is typically implemented as a probabilistic procedure using the Gumbel-max trick. However, sampling does not necessarily have to be probabilistic; it suffices that samples are generated faithfully according to the given transition probabilities. This motivates us to explore deterministic alternatives for the reverse denoising process.
@@ 2.2 Herding Algorithm
The dynamics of the herding system is classified as a piecewise isometry (Goetz, 2000; 2003), which is weakly chaotic, with Lyapunov exponents equal to zero almost everywhere, and typically has a fractal attracting set. Such complex nonlinear dynamics of the herding system yields diverse samples with relatively high entropy and negative autocorrelations (Welling, 2009; Welling & Chen, 2010).
@@
3 Derandomization of Reverse Denoising Process
We apply the herding algorithm to the reverse denoising process of discrete diffusion models. For derandomization of the Markov chain, we need to introduce a time-dependent version of the herding algorithm as in (Suzuki, 2014).
Specifically, the sample $\bm{x}_{t}$ and weight vector $\bm{w}_{t}$ for each token are updated from time $t$ to t1t-1 as follows:
@@ 3 Derandomization of Reverse Denoising Process
where $\bm{p}_{t-1}=p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$ denotes the probability vector of the categorical distribution obtained from the model prediction. In Eq. (9), the current probability vector $\bm{p}_{t-1}$ is taken into account when selecting the new sample $\bm{x}_{t-1}$, by taking the argmax of $(\bm{w}_{t}+\bm{p}_{t-1})^{\top}\bm{x}$ instead of $\bm{w}_{t}^{\top}\bm{x}$. This herding-based system reduces the discrepancy $\sum_{\tau=t}^{T-1}(\bm{p}_{\tau}-\bm{x}_{\tau})$ compared with stochastic sampling, thereby yielding a sample sequence that reflects transition probabilities more faithfully. A proof of boundedness and $O(T^{-1})$ convergence, inherited from the original herding system, is given in Section A.
@@
That is, the sample $\bm{x}_{t-1}$ remains the same as $\bm{x}_{t}$ unless another candidate exceeds the objective value by at least $\delta$.
@@
This herding-based denoising algorithm is entirely deterministic; the only source of randomness arises from the initialization of $(\bm{x}_{T},\bm{w}_{T})$, which is chosen randomly from the uniform distribution on $\mathcal{V}\times[0,1]^{K}$. By iterating the update rules, the algorithm induces a mapping from a pure noise $(\bm{x}^{(1:L)}_{T},\bm{w}^{(1:L)}_{T})$ to a sample $(\bm{x}^{(1:L)}_{0},\bm{w}^{(1:L)}_{0})$, which constitutes a piecewise isometry. Unlike the mapping defined by ODE flows in continuous diffusion models, it is not one-to-one and the probability mass is preserved piecewise. In other words, small perturbations in the weight vector persist unless they alter the argmax; however, once the generated sample changes, they cause a drastic shift in the trajectory of the dynamics, reflecting the weakly chaotic behavior with zero Lyapunov exponents almost everywhere.
@@
The time-dependent herding system introduced here can be regarded as a natural extension of the $\Delta\Sigma$-modulator (Inose et al., 1962; Inose & Yasuda, 1963) and the error diffusion algorithm (Adler et al., 2003), which have long been studied in engineering fields such as signal and image processing.
4 Experiments
We demonstrate the effectiveness of the proposed discrete denoising algorithm through experiments on text and image generation. For clarity of presentation, we reproduce the trained UDLM model as reported in (Schiff et al., 2025) and evaluate the improvements that result solely from introducing the herding algorithm into the reverse denoising process. All experiments are conducted based on the official code provided by Schiff et al. (2025).
4.1 Character-Level Text Generation
We first evaluate the proposed algorithm in character-level text generation, using the text8 dataset (Mahoney, 2011). The vocabulary size is $35$ and the output length is fixed to $256$ characters. The model architecture follows the UDLM paper (Schiff et al., 2025), which is based on Transformer with 92.4M parameters. For evaluation, we generate 1,024 samples for each condition and compute the generative perplexity based on the GPT-2 Large pretrained model (Radford et al., 2019) and the entropy.
Figure 1 shows the generative perplexity and entropy of the generated samples. The herding-based denoising with $\delta=0$ consistently outperforms stochastic denoising in terms of perplexity, achieving up to a $10\times$ improvement for $T=1024$ inference steps, albeit at the expense of entropy. Increasing the margin $\delta$ interpolates between these two extremes, allowing us to improve perplexity with a moderate trade-off in entropy.
Figure 1: Evaluation of text8 character-level text samples by generative perplexity (Gen. PPL) and entropy. Results for stochastic denoising (Gumbel-max) and herding-based denoising with different $\delta$ values are shown.
4.2 Word-Level Text Generation
We also evaluate the proposed algorithm in word-level text generation, using the one billion words dataset (LM1B) (Chelba et al., 2013). The vocabulary size is $30,522$ and the output length is $128$ tokens. The model architecture follows the UDLM paper (Schiff et al., 2025), which is based on Transformer with 139M parameters. For evaluation, we generate 1,024 samples for each condition and compute the generative perplexity and MAUVE score (Pillutla et al., 2021) based on the GPT-2 Large pretrained model (Radford et al., 2019) as well as the entropy.
Table 1 shows the evaluation results. In our reproduction experiment, the UDLM model trained using wrapped LM1B sequences performs worse than the published results (Schiff et al., 2025). Nevertheless, the herding-based denoising substantially reduces the perplexity compared to stochastic denoising for both $T=1024$ and $T=128$ steps, outperforming the published UDLM (uniform) and MDLM (masked) results and even the autoregressive (AR) model. We also observe a slight loss in entropy and a small improvement in the MAUVE score. Overall, the herding-based denoising significantly improves the coherence and linguistic quality of the generated text while affecting the distributional metrics only minimally. It should also be noted that the generative perplexity improves as the number of inference steps $T$ increases, which suggests that iterative refinement is enhanced by deterministic denoising.

### v3

@@ Abstract
We propose a deterministic denoising algorithm for discrete-state diffusion models. The key idea is to derandomize the generative reverse Markov chain by introducing a variant of the herding algorithm, which induces deterministic state transitions driven by weakly chaotic dynamics. It serves as a direct replacement for the stochastic denoising process, without requiring retraining or continuous state embeddings. We demonstrate consistent improvements in both efficiency and sample quality on text and image generation tasks. In addition, the proposed algorithm yields improved solutions for diffusion-based combinatorial optimization. Thus, herding-based denoising is a simple yet promising approach for enhancing the generative process of discrete diffusion models. Furthermore, our results reveal that deterministic reverse processes, well established in continuous diffusion, can also be effective in discrete state spaces.
discrete diffusion, deterministic denoising, herding algorithm, discrete generative modeling, combinatorial optimization
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have recently achieved remarkable success in generating realistic image and audio data. Their forward process progressively corrupts data with noise, and the models are trained to denoise by inverting the corruption process. Once trained, diffusion models can generate realistic samples from pure noise by iteratively applying the denoising procedure. For continuous data such as images and speech, the reverse denoising process is often implemented as a deterministic algorithm, which exhibits more directed dynamics and enables efficient generation with fewer denoising steps compared to stochastic denoising, as demonstrated by DDIM (Song et al., 2021).
@@
Diffusion models for discrete data (Hoogeboom et al., 2021; Austin et al., 2021) are also an important research direction, since many data of interest are inherently discrete, including text, graphs, genomes, and chemical structures. In discrete-state diffusion models, the forward and reverse processes are typically formulated as probabilistic Markov chains on a discrete sample space (Austin et al., 2021). If the reverse denoising process could be made deterministic, improved performance would be expected, as in the case of continuous models. However, in discrete spaces, naive derandomization as a deterministic mapping aggregates discrete trajectories and reduces sample diversity, which is essential for generative modeling.
@@
One promising approach is to embed the discrete sample space into a continuous space and apply continuous diffusion models (Hoogeboom et al., 2021; Chen et al., 2023; Sahoo et al., 2025). The main advantage of this approach is that it allows the direct use of well-established techniques for continuous diffusion, including deterministic denoising. Nevertheless, denoising in the continuous space requires retraining the model. Moreover, it remains unclear whether embedding discrete data into a continuous space is truly necessary for derandomization.
@@
In this paper, we propose to directly derandomize the reverse denoising process on discrete sample spaces. The key idea is to introduce a variant of the herding algorithm (Welling, 2009; Welling & Chen, 2010), which associates auxiliary continuous weight variables with each discrete variable in the sample space. The reverse denoising updates of both discrete and continuous variables are deterministic, and randomness arises solely from their initial states. Since the proposed algorithm derandomizes the state transitions of Markov chains, it serves as a drop-in replacement for the stochastic reverse process without retraining discrete diffusion models. This relation is analogous to DDIM (Song et al., 2021), which generates samples deterministically using a model trained for DDPM (Ho et al., 2020), as summarized in Table 1. In our experiments on text and image generation, we demonstrate consistent improvements in both efficiency and sample quality. These results reveal that deterministic reverse processes, well established in continuous diffusion, can also be effective in discrete state spaces.
Table 1: Deterministic denoising algorithms for diffusion models with continuous, masked discrete, and uniform discrete noising processes. We establish a herding-based deterministic denoising algorithm for discrete diffusion models without masking.
Stochastic Deterministic
Continuous DDPM (Ho et al., 2020) DDIM (Song et al., 2021)
Discrete Masked D3PM absorbing (Austin et al., 2021) MDLM (Sahoo et al., 2024) DNDM (Chen et al., 2024)
Discrete Uniform D3PM uniform (Austin et al., 2021) UDLM (Schiff et al., 2025) herding-based denoising (ours)
@@
For evaluation, we focus on discrete uniform diffusion models, while the proposed denoising algorithm is applicable to various noising processes. Specifically, in our experiments, we adopt UDLM (Schiff et al., 2025), a state-of-the-art discrete-state generative model based on the uniform diffusion process. Our aim in this paper is not to propose a new discrete diffusion model, but to demonstrate how the herding-based algorithm improves the generative process of existing models. We use the official training, sampling, and evaluation codes of UDLM, modifying only about 20 lines in the reverse denoising procedure. This directly demonstrates that our method works as a drop-in replacement without retraining. Moreover, the uniform diffusion process is particularly suitable for our approach, since its reverse process performs iterative refinement through successive Markov state transitions, making it a natural setting to examine whether derandomizing these transitions can improve efficiency and sample quality. As another evaluation axis, we apply the proposed algorithm to DIFUSCO (Sun & Yang, 2023), a diffusion-based combinatorial optimization framework, to assess improvements in objective values achieved by the derandomized reverse process.
Diffusion language models are an important research area within discrete diffusion, while our approach is not restricted to text data. In this context, masked diffusion models have been extensively studied and are expected to enable parallel and coherent text generation with iterative refinement. While autoregressive (AR) models (Bengio et al., 2003) generate tokens sequentially, masked diffusion models allow tokens to be generated in arbitrary orders, which can offer flexibility in generation. Early masked diffusion models can be regarded as relatively static in their reverse processes, in the sense that each token transitions only once from the masked state and remains fixed thereafter, which is a consequence of the absorbing noising process toward the all-mask state. As summarized in Table 1, derandomization for such static masked diffusion has been explored in DNDM (Chen et al., 2024), where the reverse process is derandomized by externally specifying the unmasking transition times. More recent studies, however, have moved toward more dynamic reverse processes that enable iterative refinement, where tokens may be remasked, as in LLaDA (Nie et al., 2025; Zhu et al., 2025), or resampled multiple times, as in GIDD (von Rütte et al., 2025). The uniform diffusion process can be considered the extreme case of this trend, where Markov state transitions in the token space continue throughout the reverse process (Schiff et al., 2025; von Rütte et al., 2025). Notably, recent work suggests that uniform diffusion scales more efficiently than masked diffusion in data-limited large-parameter regimes (von Rütte et al., 2025), which is particularly relevant given the growing concern over data scarcity (Villalobos et al., 2024). Accordingly, in diffusion-based text generation, our derandomization approach is expected to facilitate effective iterative refinement in dynamic reverse processes driven by Markov chains.
Overall, we present a simple yet effective derandomization approach that directly enhances the generative process of existing discrete-state diffusion models. The proposed algorithm exhibits weakly chaotic, or piecewise isometric dynamics, which preserves distance and probability mass locally, with a superior asymptotic convergence rate compared with stochastic sampling. Thus, this work provides a new perspective for viewing discrete diffusion models through the notions of nonlinear dynamics and deterministic sampling.
2 Preliminaries
@@ 2.1 Discrete Diffusion Models
A diffusion process in a discrete state space can be viewed as a sequence of discrete state transitions that gradually corrupt the original data. This process can be formulated as a discrete-time Markov chain over categorical variables, giving rise to a reverse-time denoising Markov chain (Austin et al., 2021). This formulation can be naturally extended to state transitions in continuous time, using Markov jump processes parameterized by transition-rate matrices (Campbell et al., 2022; Sun et al., 2023). Foundational theoretical studies analyze approximation error and convergence properties of both discrete-time and continuous-time discrete diffusion models (Ren et al., 2025a; Zhang et al., 2025; Liang et al., 2025a, b). Below, we focus on a discrete-time formulation to outline how our approach can be applied to the reverse denoising process.
We consider generative models for discrete data represented as a sequence of $L$ categorical variables. Each variable takes one of $K$ categories and is represented as a $K$-dimensional one-hot vector $\bm{x}\in\mathcal{V}$, where $\mathcal{V}=\{(x_{1},x_{2},\dots,x_{K})^{\top}\in\{0,1\}^{K}\mid\sum_{k=1}^{K}x_{k}=1\}$. A sequence of $L$ variables is denoted by $\bm{x}^{(1:L)}=(\bm{x}^{(1)},\bm{x}^{(2)},\dots,\bm{x}^{(L)})\in\mathcal{V}^{L}$.
@@
Algorithm 1 Stochastic denoising for uniform diffusion
0: trained prediction model $p_{\theta}$
1:for $l=1$ to $L$ do
2:  $\bm{x}^{(l)}_{T}\sim\mathrm{Uniform}(\mathcal{V})$
3:end for
4:for $t=T$ down to 11 do
5:  for $l=1$ to $L$ do
6:   $\bm{p}^{(l)}_{t-1}\leftarrow p_{\theta}(\bm{x}^{(l)}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$
7:   $\bm{x}^{(l)}_{t-1}\sim\mathrm{Cat}(\bm{x}^{(l)}_{t-1};\bm{p}^{(l)}_{t-1})$
8:  end for
9:end for
10:Output: generated sample $\bm{x}^{(1:L)}_{0}$
By iterating this Markov transition from pure noise $\bm{x}^{(1:L)}_{T}$ sampled from the stationary distribution, the reverse process generates a sample $\bm{x}^{(1:L)}_{0}$, as described in Algorithm 1. Namely, each denoising step reduces to sampling from the categorical distribution $\mathrm{Cat}(\bm{x}_{t-1};\bm{p}_{t-1})$ with probability vector $\bm{p}_{t-1}=p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$. This is typically implemented as a probabilistic procedure using the Gumbel-max trick. However, sampling does not necessarily have to be probabilistic; it suffices that samples are generated faithfully according to the given transition probabilities. This motivates us to explore deterministic alternatives for the reverse denoising process.
@@ 2.2 Herding Algorithm
The dynamics of the herding system is classified as a piecewise isometry (Goetz, 2000, 2003), which is weakly chaotic, with Lyapunov exponents equal to zero almost everywhere, and typically has a fractal attracting set. Such complex nonlinear dynamics of the herding system yields diverse samples with relatively high entropy and negative autocorrelations (Welling, 2009; Welling & Chen, 2010).
@@
3 Herding-Based Denoising
3.1 Derandomization of Reverse Markov Chain
We apply the herding algorithm to derandomize the reverse Markov chain of discrete-state diffusion models. Since the transition probability at each denoising step is time-dependent, we need to introduce a time-dependent version of the herding algorithm as in (Suzuki, 2014).
Specifically, the sample $\bm{x}_{t}$ and weight vector $\bm{w}_{t}$ are updated from time $t$ to t1t-1 as follows:
@@ 3.1 Derandomization of Reverse Markov Chain
where $\bm{p}_{t-1}=p_{\theta}(\bm{x}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$ denotes the probability vector of the categorical distribution obtained from the model prediction. Since the transition probability is time-dependent, the current probability vector $\bm{p}_{t-1}$ is taken into account when selecting the new sample $\bm{x}_{t-1}$ in Eq. (9), as opposed to the original herding update in Eq. (7).
Algorithm 2 Herding-based denoising
0: trained prediction model $p_{\theta}$
1:for $l=1$ to $L$ do
2:  $\bm{x}^{(l)}_{T}\sim\mathrm{Uniform}(\mathcal{V})$
3:  $\bm{w}^{(l)}_{T}\sim\mathrm{Uniform}([0,1]^{K})$
4:end for
5:for $t=T$ down to 11 do
6:  for $l=1$ to $L$ do
7:   $\bm{p}^{(l)}_{t-1}\leftarrow p_{\theta}(\bm{x}^{(l)}_{t-1}\mid\bm{x}^{(1:L)}_{t},t)$
8:   $\bm{x}^{(l)}_{t-1}\leftarrow\operatorname*{arg\,max}_{\bm{x}\in\mathcal{V}}(\bm{w}^{(l)}_{t}+\bm{p}^{(l)}_{t-1}+\delta\bm{x}^{(l)}_{t})^{\top}\bm{x}$
9:   $\bm{w}^{(l)}_{t-1}\leftarrow\bm{w}^{(l)}_{t}+\bm{p}^{(l)}_{t-1}-\bm{x}^{(l)}_{t-1}$
10:  end for
11:end for
12:Output: generated sample $\bm{x}^{(1:L)}_{0}$
Figure 1: Evaluation of text8 character-level text samples by generative PPL (Gen. PPL) and entropy. Results for stochastic denoising (Gumbel-max) and herding-based denoising with different $\delta$ values are shown.
@@
That is, the sample $\bm{x}_{t-1}$ remains the same as $\bm{x}_{t}$ unless another candidate exceeds the objective value by at least $\delta$. A pseudocode of this algorithm is provided in Algorithm 2.
This herding-based system reduces the cumulative discrepancy $\sum_{\tau=t}^{T-1}(\bm{p}_{\tau}-\bm{x}_{\tau})$ compared with stochastic sampling, thereby yielding a sample sequence that reflects transition probabilities more faithfully. The boundedness and $O(T^{-1})$ discrepancy bound, inherited from the original herding system, are proved in Section 3.2.
@@
As the proposed denoising algorithm is entirely deterministic, the only source of randomness arises from the initialization of $(\bm{x}_{T},\bm{w}_{T})$, which is drawn randomly from the uniform distribution on $\mathcal{V}\times[0,1]^{K}$. Through the iterative updates, the algorithm induces a mapping from a pure noise $(\bm{x}^{(1:L)}_{T},\bm{w}^{(1:L)}_{T})$ to a sample $(\bm{x}^{(1:L)}_{0},\bm{w}^{(1:L)}_{0})$, which constitutes a piecewise isometry. Unlike the mapping defined by ODE flows in continuous diffusion models, this mapping is not one-to-one, and the probability mass is preserved piecewise. In other words, small perturbations in the weight vector persist unless they alter the argmax. Nevertheless, once the generated sample changes, these perturbations cause a drastic shift in the trajectory, reflecting the weakly chaotic dynamics.
@@
The time-dependent herding system introduced here converts a $K$-dimensional real-valued sequence $\{\bm{p}_{t}\}$ into a one-hot discrete sequence $\{\bm{x}_{t}\}$. In this sense, it can also be viewed as a multidimensional extension of 1-bit analog-to-digital (A/D) converters that represent continuous signals using discrete binary signals, as exemplified by the $\Delta\Sigma$-modulator (Inose et al., 1962; Inose & Yasuda, 1963) and the error diffusion algorithm (Adler et al., 2003), which have long been studied in engineering fields such as signal and image processing.
3.2 Bounded Dynamics, Convergence, and Initialization
As a theoretical justification for our proposed algorithm, we provide a proof of the boundedness of the weight vector $\bm{w}_{t}$ and the $O(T^{-1})$ bound on the discrepancy between the empirical sample distribution and the averaged probability vector.
Proposition 1.
For any initial weight $\bm{w}_{T}\in\mathbb{R}^{K}$, the infinite reverse orbit $\{\bm{w}_{t}\}_{t\leq T}$ generated by the herding-based denoising is bounded.
$\sum_{i}w_{t-1,i}$ $=\sum_{i}(w_{t,i}+p_{t-1,i}-x_{t-1,i})$
$=\sum_{i}w_{t,i}+1-1=\sum_{i}w_{t,i}.$
Let $\mu=\sum_{i}w_{t,i}/K$ be the average weight. We introduce a constant $\ell=\min\{\mu-\delta-1,\min_{i}w_{T,i}\}$, which satisfies $w_{T,i}\geq\ell$ for all $i$ at the initial time $T$. Now assume that $w_{t,i}\geq\ell$ holds for some $t$. If $x_{t-1,i}=0$, then $w_{t-1,i}=w_{t,i}+p_{t-1,i}-x_{t-1,i}\geq w_{t,i}\geq\ell$. If $x_{t-1,i}=1$, then $w_{t-1,i}=w_{t,i}+p_{t-1,i}-x_{t-1,i}\geq\max_{j}(w_{t,j}+p_{t-1,j})-\delta-1\geq\mu-\delta-1\geq\ell$. Thus, in both cases, we obtain $w_{t-1,i}\geq\ell$, and by induction the inequality holds for all $t$. This establishes a uniform lower bound $\ell$ for each component of $\bm{w}_{t}$. Since a lower bound exists and the total weight $\sum_{i}w_{t,i}$ is conserved, every component of $\bm{w}_{t}$ must remain bounded from above as well, and therefore $\bm{w}_{t}$ is bounded. ∎
Proposition 2.
The discrepancy between the averaged probability vector $\langle\bm{p}_{\tau}\rangle_{\tau=0}^{T-1}$ and the empirical distribution $\langle\bm{x}_{\tau}\rangle_{\tau=0}^{T-1}$ of samples generated by the herding-based denoising converges to zero as $T\to\infty$ with convergence rate $O(T^{-1})$, where $\langle\cdot\rangle_{\tau=0}^{T-1}$ denotes the average over $\tau=0,1,\dots,T-1$.
From Eq. (10), the discrepancy can be rewritten as
$\langle\bm{x}_{\tau}-\bm{p}_{\tau}\rangle_{\tau=0}^{T-1}=\langle\bm{w}_{\tau+1}-\bm{w}_{\tau}\rangle_{\tau=0}^{T-1}=\frac{1}{T}(\bm{w}_{T}-\bm{w}_{0}),$
which converges to $\bm{0}$ with rate $O(T^{-1})$ as $T\to\infty$, due to the boundedness of $\bm{w}_{t}$. ∎
Although the asymptotic convergence is guaranteed for any initial weight vector $\bm{w}_{T}\in\mathbb{R}^{K}$, its initialization plays an important role in practical sample generation for finite $T$.
For uniform diffusion noising, random initialization of $\bm{w}_{T}$ from $0,1]^{K}$ is justified for the following reason. The stationary distribution of the uniform diffusion Markov chain is the uniform distribution over the sample space $\mathcal{V}$. Therefore, we regard $\bm{w}_{T}$ as a weight vector obtained after iterating the herding updates infinitely many times with the uniform probability vector fixed at $\bm{p}_{T}=\bm{1}/K$. After a transient period, each weight $w_{t,i}$ increases by $1/K$ at each iteration but decreases by $1$ once every $K$ iterations. As a result, all $w_{T,i}$ remain within an interval of length $1$. Accordingly, we choose the uniform distribution on $0,1]^{K}$ as a natural initialization.